# FASTopoWM: Fast-Slow Lane Segment Topology Reasoning with Latent World Models

### v1

@@ Abstract
Lane segment topology reasoning provides comprehensive bird’s-eye view (BEV) road scene understanding, which can serve as a key perception module in planning-oriented end-to-end autonomous driving systems. Existing lane topology reasoning methods often fall short in effectively leveraging temporal information to enhance detection and reasoning performance. Recently, stream-based temporal propagation method has demonstrated promising results by incorporating temporal cues at both the query and BEV levels. However, it remains limited by over-reliance on historical queries, vulnerability to pose estimation failures, and insufficient temporal propagation. To overcome these limitations, we propose FASTopoWM, a novel fast-slow lane segment topology reasoning framework augmented with latent world models. To reduce the impact of pose estimation failures, this unified framework enables parallel supervision of both historical and newly initialized queries, facilitating mutual reinforcement between the fast and slow systems. Furthermore, we introduce latent query and BEV world models conditioned on the action latent to propagate the state representations from past observations to the current timestep. This design substantially improves the performance of temporal perception within the slow pipeline. Extensive experiments on the OpenLane-V2 benchmark demonstrate that FASTopoWM outperforms state-of-the-art methods in both lane segment detection (37.4% v.s. 33.6% on mAP) and centerline perception (46.3% v.s. 41.5% on OLS).
Code — https://github.com/YimingYang23/FASTopoWM
Introduction
@@
Lane segment topology reasoning predicts lane segments (including centerlines and boundary lines) along with their topological relationships to construct a comprehensive road network (Wang et al. 2024a; Li et al. 2023b). This capability can be integrated into planning-oriented end-to-end autonomous driving systems, serving as the perception module to provide bird’s-eye view (BEV) road scene understanding (Hu et al. 2023b; Jiang et al. 2023; Zhou et al. 2025).
@@
Most existing approaches (Liao et al. 2022; Fu et al. 2025a) focus on single-frame detection, failing to exploit the temporal consistency of predictions across consecutive frames. This limitation leads to inconsistent topology and position of lane segments in the global coordinate system. To overcome these issues, stream-based temporal propagation methods (Yuan et al. 2024) have proven effective in enhancing temporal consistency in perception. They reuse high-confidence historical queries and BEV features as anchors for current detection. Inspired by this observation, we introduce the stream-based approach in this work to predict lane segment topology with temporal consistency. However, as illustrated in Fig. 1, current stream-based frameworks suffer from three critical limitations: (1) Over-reliance on historical queries. The historical queries demonstrate higher confidence levels, they are more likely to approximate the GT positions compared to the newly initialized queries. As a result, during Hungarian matching supervision, historical queries are prioritized while newly initialized queries are often neglected. However, in the first frame of a scene, only the newly initialized queries are available for lane segment detection. If their performance is suboptimal, errors may propagate and accumulate across subsequent frames. (2) Vulnerability to pose estimation failures. To align historical queries and BEV features with the current frame, stream-based approaches rely on continuous pose estimation between adjacent frames. However, when vehicles enter tunnels or remote areas where GPS signals are unavailable, or when IMU error accumulation prevents pose refinement, the system degrades to single-frame detection mode. This degradation leads to significant performance deterioration and may ultimately cause complete system failure. (3) Weak temporal propagation. When historical BEV features are warped, the details of the features at the edges of the BEV tend to be lost. Moreover, simply applying pose transformation to historical queries through a basic MLP architecture fails to achieve optimal performance.
@@
To address the aforementioned issues, we propose a novel fast-slow lane segment topology reasoning framework with latent world models (FASTopoWM). Inspired by recent vision-language models (VLMs) (Zhang et al. 2025; Xiao et al. 2025), we decouple our network into dual pathways: a slow pipeline and a fast pipeline. The slow pipeline leverages temporal information to address challenging perception scenarios and improve detection performance, while the fast pipeline performs single-frame perception to ensure the basic functionality of the system. The two systems can perform inference in parallel or operate independently. A key innovation lies in the unified framework between the fast and slow systems that allowing parallel supervision of both historical and initialized queries. This eliminates the need for training separate models for temporal and single-frame detection and enables mutual reinforcement between the fast and slow pipelines. Specifically, the slow pipeline benefits from a better initialization for temporal propagation, while the fast pipeline implicitly leverages the temporal dynamics learned through shared weights. To improve temporal propagation, we propose latent query and BEV world models based on the principle that ”the present represents a continuation of the past.” Conditioned on the relative poses in adjacent frames, both latent world models propagate state representations from historical observations to the current timestep, significantly enhancing the robustness of temporal propagation in the slow pipeline. Our contributions are summarized as follows:
We identify severe performance degradation issues in existing stream-based methods. To address this, we propose FASTopoWM, a novel fast-slow framework augmented with latent world models for robust lane segment topology reasoning.
@@
We introduce a unified fast-slow system that enables parallel supervision of both historical and newly initialized queries. This design facilitates mutual reinforcement and allows inference switching based on pose estimation conditions, thereby enhancing system robustness.
We design two latent world models that effectively capture temporal dynamics and enable strong temporal propagation.
FASTopoWM is evaluated on the OpenLane-V2 dataset (Wang et al. 2024a), achieving state-of-the-art performance in lane topology reasoning.
Related Work
HD Map and Lane Topology Reasoning
@@
Current high-definition (HD) map learning frameworks employ end-to-end detection pipelines to generate vectorized representations of map elements. VectorMapNet (Liu et al. 2023) formulates map elements as polyline to eliminate heuristic post-processing. MapTR (Liao et al. 2022, 2023) develops permutation-equivalent modeling for lane point sets. Mask2map (Choi et al. 2024) incorporates mask-aware queries and BEV features to enhance semantic understanding. MapDR (Chang et al. 2025) decomposes road networks into geometric, connectivity, and regulatory layers. For temporal modeling, StreamMapNet (Yuan et al. 2024) implements stream-based propagation. SQD-MapNet (Wang et al. 2024b) applies query denoising for BEV boundary consistency. InteractionMap (Wu, Yang, and Li 2025) achieves comprehensive temporal fusion through local-to-global integration. In contrast to HD map learning, lane topology reasoning primarily focuses on the topological relationships among lanes. TopoNet (Li et al. 2023a) establishes dual lane-to-lane and lane-to-traffic graphs. TopoLogic (Fu et al. 2025a) and TopoPoint (Fu et al. 2025b) emphasize the critical role of endpoints in topological connectivity. TopoMLP (Wu et al. 2023) explicitly encodes coordinate information to improve topological reasoning. Topo2Mask (Kalfaoglu et al. 2024) introduces instance masks and mask attention mechanisms to update query representations. TopoFormer (Lv et al. 2025) proposes geometry-guided and counterfactual self-attention to enhance road topology understanding. In contrast to these approaches, we propose a fast-slow system that addresses the limitations of existing stream-based methods and incorporate two latent world models to effectively capture temporal dynamics.
World Model in Autonomous Driving
@@
The world model serves as a bridge between understanding and generating future states. In autonomous driving, it can predict the future state of the ego vehicle conditioned on its actions (Bar et al. 2025). GAIA-1 (Hu et al. 2023a) leverages sequence learning to anticipate future events. DriveDreamer (Wang et al. 2024c) introduces a two-stage diffusion training pipeline to generate controllable driving scenes. Instead of synthesizing pixel-level content, OccWorld (Zheng et al. 2024) forecasts 3D occupancy to simulate future scenarios. To reduce computational overhead and modeling complexity, recent approaches have shifted toward predicting compact latent representations using world models. BEVWorld (Zhang et al. 2024) proposes a latent BEV sequence diffusion model to forecast future scenes conditioned on multi-modal inputs. WoTE (Li et al. 2025) predicts future BEV features to evaluate high-confidence trajectories. LAW (Li et al. 2024) employs a latent world model for self-supervised learning. Motivated by prior work, we design a dual latent world model framework that learns the transformation from past to present, enabling more effective temporal reasoning and enhancing current lane topology prediction.
Fast-Slow System
@@
In autonomous driving scenarios, where real-time performance and reliability are essential, adding computational burden or auxiliary cues to boost prediction accuracy is often infeasible in practice. As a trade-off solution, fast-slow systems have recently been introduced. The fast system performs essential predictions with minimal overhead, while the slow system leverages additional computation and auxiliary cues to handle complex scenarios. Fast-slow systems enable flexible mode switching tailored to the varying demands of different scenarios. Chameleon (Zhang et al. 2025) utilizes a VLM with chain-of-thought (CoT) reasoning to conduct neuro-symbolic topology inference in the slow system. FAST (Xiao et al. 2025) dynamically switches between short and long reasoning paths based on task complexity. SlowFast-LLaVA (Xu et al. 2024) employs different sampling rates to capture motion cues and fuses slow and fast features to efficiently represent video information. Xu et al. (Xu et al. 2025) combine a “slow” LLM for command parsing with a “fast” RL agent for vehicle control. FASIONAD++ (Qian et al. 2025) incorporates a VLM as the slow system to provide feedback and evaluation for a fast end-to-end pipeline. However, in many of these approaches, the fast and slow systems are implemented as entirely separate frameworks, and switching occurs based on the scenario. In contrast, our method does not require additional models for the fast or slow pathways. Instead, fast and slow systems are trained under parallel supervision, enabling mutual reinforcement and improving overall robustness.
Method
Problem Formulation
@@
Overview
Fig. 2 illustrates the framework of FASTopoWM. The overall architecture can be briefly divided into encoder and decoder parts. Both the slow and fast systems are integrated into a unified network. The PV-to-BEV encoder (Li et al. 2022; He et al. 2016; Lin et al. 2017) extracts BEV features , where C, H, W represent the number of feature channels, height, and width, respectively. The initialized queries are derived from a learnable embedding space. is the number of queries. The relative ego poses between adjacent frames are encoded into an action latent . Conditioned on this action latent, the query world model and the BEV world model transform the historical queries and BEV features , respectively, to generate the stream queries and stream BEV features for the current frame. The stream BEV features are then fused with the BEV features extracted from the current frame. The first transformer layer takes the initialized queries and the BEV features from the current frame as input. The subsequent transformer layers share weights and receive parallel inputs from both the slow and fast systems. The slow system incorporates temporal information, whereas the fast system does not. During training, both systems are supervised jointly. At inference time, the prediction from the slow system is used as the final output when reliable pose information is available. Otherwise, when such information is missing or inaccurate, the prediction from the fast system is adopted. This design improves robustness while leveraging temporal cues to enhance performance when conditions permit.
Temporal Propagation via Latent World Models
Temporal propagation leverages detection results from previous frames as auxiliary information to assist prediction in the subsequent frame (Yuan et al. 2024). In contrast to existing latent world model methods (Li et al. 2024), our approach eliminates the need for future trajectory prediction. Instead, grounded in the principle that “the present represents a continuation of the past,” we utilize known relative poses to infer the current state based on past observations.
@@
where query world model is composed of Transformer blocks, containing self-attention modules and feed-forward modules. Similarly, the BEV world model comprises Transformer blocks that incorporate temporal self-attention modules (Li et al. 2022) and feed-forward modules.
Then, the stream BEV feature are fused with extracted BEV feature using gated recurrent unit (Chung et al. 2014) to enrich temporal cues.
Future Latent Supervision. Unlike previous methods that rely on warping and often lose information at the BEV boundary (Wang et al. 2024b), our BEV world model imagines the evolution of the next-frame BEV representation based on historical BEV features and relative poses. Since dense BEV annotations are difficult to obtain, we adopt a self-supervised learning strategy based on BEV features from temporally adjacent frames. We use mean squared error (MSE) loss to align the stream BEV features with the extracted BEV features of the current frame:
For queries, we employ transformation loss (Yang et al. 2025b) to supervise consistency of coordinate, category, and semantic BEV mask:
where , , and are coordinates of lane segment, classes of centerline, boundary types and semantic BEV mask predicted from stream queries . , , and are GT annotations transformed from T-1 frame to T frame. The overall future latent supervision can be expressed as:
@@
| Method | Epochs | Temporal | mAP | FPS | ||||
| MapTR (Liao et al. 2022) | 24 | No | 27.0 | 25.9 | 28.1 | - | - | 14.5 |
| MapTRv2 (Liao et al. 2023) | 24 | No | 28.5 | 26.6 | 30.4 | - | - | 13.6 |
| TopoNet (Li et al. 2023a) | 24 | No | 23.0 | 23.9 | 22.0 | - | - | 10.5 |
| LaneSegNet (Li et al. 2023b) | 24 | No | 32.6 | 32.3 | 32.9 | 25.4 | 45.9 | 14.7 |
| TopoLogic (Fu et al. 2025a) | 24 | No | 33.2 | 33.0 | 33.4 | 30.8 | - | - |
| Topo2Seq (Yang et al. 2025a) | 24 | No | 33.6 | 33.7 | 33.5 | 26.9 | 48.1 | 14.7 |
| FASTopoWM (ours) | 24 | No | 34.1 | 33.9 | 34.4 | 26.6 | 48.2 | 14.0 |
| FASTopoWM† (ours) | 24 | No | 34.2 | 34.0 | 34.4 | 28.4 | 48.2 | 14.0 |
| StreamMapNet (Yuan et al. 2024) | 24 | Yes | 20.3 | 22.1 | 18.6 | 13.2 | 33.2 | 14.1 |
| SQD-MapNet (Wang et al. 2024b) | 24 | Yes | 26.0 | 27.1 | 24.9 | 16.6 | 39.4 | 14.1 |
| FASTopoWM (ours) | 24 | Yes | 37.4 | 36.4 | 38.4 | 29.6 | 51.2 | 11.4 |
| FASTopoWM† (ours) | 24 | Yes | 37.2 | 36.2 | 38.1 | 31.6 | 51.3 | 11.4 |
| Method | Epochs | Temporal | |||
| VectorMapNet (Liu et al. 2023) | 24 | No | - | 3.5 | - |
| STSU (Can et al. 2021) | 24 | No | - | 8.2 | - |
| MapTR (Liao et al. 2022) | 24 | No | - | 15.2 | - |
| TopoNet (Li et al. 2023a) | 24 | No | 25.1 | 24.3 | 6.7 |
| TopoMLP (Wu et al. 2023) | 24 | No | 36.2 | 26.6 | 19.8 |
| LaneSegNet (Li et al. 2023b) | 24 | No | 38.7 | 27.5 | 24.9 |
| TopoLogic (Fu et al. 2025a) | 24 | No | 36.2 | 25.9 | 21.6 |
| TopoFormer⋆ (Lv et al. 2025) | 24 | No | 41.5 | 34.8 | 23.2 |
| FASTopoWM (ours) | 24 | No | 41.8 | 31.6 | 27.1 |
| StreamMapNet (Yuan et al. 2024) | 24 | Yes | 26.7 | 18.9 | 11.9 |
| SQD-MapNet (Wang et al. 2024b) | 24 | Yes | 29.1 | 21.9 | 13.3 |
| FASTopoWM (ours) | 24 | Yes | 46.3 | 35.1 | 33.0 |
Unified Fast-Slow Decoder

### v4

@@ Abstract
Lane segment topology reasoning provides comprehensive bird’s-eye view (BEV) road scene understanding, which can serve as a key perception module in planning-oriented end-to-end autonomous driving systems. Existing lane topology reasoning methods often fall short in effectively leveraging temporal information to enhance detection and reasoning performance. Recently, stream-based temporal propagation method has demonstrated promising results by incorporating temporal cues at both the query and BEV levels. However, it remains limited by over-reliance on historical queries, vulnerability to pose estimation failures, and insufficient temporal propagation. To overcome these limitations, we propose FASTopoWM, a novel fast-slow lane segment topology reasoning framework augmented with latent world models. To reduce the impact of pose estimation failures, this unified framework enables parallel supervision of both historical and newly initialized queries, facilitating mutual reinforcement between the fast and slow systems. Furthermore, we introduce latent query and BEV world models conditioned on the action latent to propagate the state representations from past observations to the current timestep. This design substantially improves the performance of temporal perception within the slow pipeline. Extensive experiments on the OpenLane-V2 benchmark demonstrate that FASTopoWM outperforms state-of-the-art methods in both lane segment detection (37.4% v.s. 33.6% on mAP) and centerline perception (46.3% v.s. 41.5% on OLS). Code is accessible at https://github.com/YimingYang23/FASTopoWM.
1 Introduction
@@
Lane segment topology reasoning predicts lane segments (including centerlines and boundary lines) along with their topological relationships to construct a comprehensive road network (Wang et al., 2024a; Li et al., 2023b). This capability can be integrated into planning-oriented end-to-end autonomous driving systems, serving as the perception module to provide bird’s-eye view (BEV) road scene understanding (Hu et al., 2023b; Jiang et al., 2023; Zhou et al., 2025).
@@
Most existing approaches (Liao et al., 2022; Fu et al., 2025a) focus on single-frame detection, failing to exploit the temporal consistency of predictions across consecutive frames. This limitation leads to inconsistent topology and position of lane segments in the global coordinate system. To overcome these issues, stream-based temporal propagation methods (Yuan et al., 2024) have proven effective in enhancing temporal consistency in perception. They reuse high-confidence historical queries and BEV features as anchors for current detection. Inspired by this observation, we introduce the stream-based approach in this work to predict lane segment topology with temporal consistency. However, as illustrated in Fig. 1, current stream-based frameworks suffer from three critical limitations: (1) Over-reliance on historical queries. The historical queries demonstrate higher confidence levels, they are more likely to approximate the GT positions compared to the newly initialized queries. As a result, during Hungarian matching supervision, historical queries are prioritized while newly initialized queries are often neglected. However, in the first frame of a scene, only the newly initialized queries are available for lane segment detection. If their performance is suboptimal, errors may propagate and accumulate across subsequent frames. (2) Vulnerability to pose estimation failures. To align historical queries and BEV features with the current frame, stream-based approaches rely on continuous pose estimation between adjacent frames. However, when vehicles enter tunnels or remote areas where GPS signals are unavailable, or when IMU error accumulation prevents pose refinement, the system degrades to single-frame detection mode. This degradation leads to significant performance deterioration and may ultimately cause complete system failure. (3) Weak temporal propagation. When historical BEV features are warped, the details of the features at the edges of the BEV tend to be lost. Moreover, simply applying pose transformation to historical queries through a basic MLP architecture fails to achieve optimal performance.
@@
To address the aforementioned issues, we propose a novel fast-slow lane segment topology reasoning framework with latent world models (FASTopoWM). Inspired by recent vision-language models (VLMs) (Zhang et al., 2025; Xiao et al., 2025), we decouple our network into dual pathways: a slow pipeline and a fast pipeline. The slow pipeline leverages temporal information to address challenging perception scenarios and improve detection performance, while the fast pipeline performs single-frame perception to ensure the basic functionality of the system. These two systems can perform inference in parallel or operate independently. A key innovation lies in the unified framework between the fast and slow systems that allowing parallel supervision of both historical and initialized queries. This eliminates the need for training separate models for temporal and single-frame detection and enables mutual reinforcement between the fast and slow pipelines. Specifically, the slow pipeline benefits from a better initialization for temporal propagation, while the fast pipeline implicitly leverages the temporal dynamics learned through shared weights. To improve temporal propagation, we propose latent query and BEV world models based on the principle that ”the present represents a continuation of the past.” Conditioned on the relative poses in adjacent frames, both latent world models propagate state representations from historical observations to the current timestep, significantly enhancing the robustness of temporal propagation in the slow pipeline.
@@
Contributions: (1) We identify severe performance degradation issues in existing stream-based methods. To address this, we propose FASTopoWM, a novel fast-slow framework augmented with latent world models for robust lane segment topology reasoning. (2) We introduce a unified fast-slow system that enables parallel supervision of both historical and newly initialized queries. This design facilitates mutual reinforcement and allows inference switching based on pose estimation conditions, thereby enhancing system robustness. (3) We design two latent world models that effectively capture temporal dynamics and enable strong temporal propagation. (4) FASTopoWM is evaluated on the OpenLane-V2 dataset (Wang et al., 2024a), achieving state-of-the-art performance in lane topology reasoning.
2 Related Work
2.1 HD Map and Lane Topology Reasoning
@@
Current high-definition (HD) map learning frameworks employ end-to-end detection pipelines to generate vectorized representations of map elements. VectorMapNet (Liu et al., 2023) formulates map elements as polyline to eliminate heuristic post-processing. MapTR (Liao et al., 2022; 2023) develops permutation-equivalent modeling for lane point sets. Mask2map (Choi et al., 2024) incorporates mask-aware queries and BEV features to enhance semantic understanding. MapDR (Chang et al., 2025) decomposes road networks into geometric, connectivity, and regulatory layers. For temporal modeling, StreamMapNet (Yuan et al., 2024) implements stream-based propagation. SQD-MapNet (Wang et al., 2024b) applies query denoising for BEV boundary consistency. InteractionMap (Wu et al., 2025) achieves comprehensive temporal fusion through local-to-global integration. In contrast to HD map learning, lane topology reasoning primarily focuses on the topological relationships among lanes. TopoNet (Li et al., 2023a) establishes dual lane-to-lane and lane-to-traffic graphs. TopoLogic (Fu et al., 2025a) and TopoPoint (Fu et al., 2025b) emphasize the critical role of endpoints in topological connectivity. TopoMLP (Wu et al., 2023) explicitly encodes coordinate information to improve topological reasoning. Topo2Mask (Kalfaoglu et al., 2024) introduces instance masks and mask attention mechanisms to update query representations. TopoFormer (Lv et al., 2025) proposes geometry-guided and counterfactual self-attention to enhance road topology understanding. In contrast to these approaches, we propose a fast-slow system that addresses the limitations of existing stream-based methods and incorporate two latent world models to effectively capture temporal dynamics.
2.2 World Model in Autonomous Driving
@@
The world model serves as a bridge between understanding and generating future states. In autonomous driving, it can predict the future state of the ego vehicle conditioned on its actions (Bar et al., 2025). GAIA-1 (Hu et al., 2023a) leverages sequence learning to anticipate future events. DriveDreamer (Wang et al., 2024c) introduces a two-stage diffusion training pipeline to generate controllable driving scenes. Instead of synthesizing pixel-level content, OccWorld (Zheng et al., 2024) forecasts 3D occupancy to simulate future scenarios. To reduce computational overhead and modeling complexity, recent approaches have shifted toward predicting compact latent representations using world models. BEVWorld (Zhang et al., 2024) proposes a latent BEV sequence diffusion model to forecast future scenes conditioned on multi-modal inputs. WoTE (Li et al., 2025) predicts future BEV features to evaluate high-confidence trajectories. LAW (Li et al., 2024) employs a latent world model for self-supervised learning. Motivated by prior work, we design a dual latent world model framework that learns the transformation from past to present, enabling more effective temporal reasoning and enhancing current lane topology prediction.
2.3 Fast-Slow System
@@
In autonomous driving scenarios, where real-time performance and reliability are essential, adding computational burden or auxiliary cues to boost prediction accuracy is often infeasible in practice. As a trade-off solution, fast-slow systems have recently been introduced. The fast system performs essential predictions with minimal overhead, while the slow system leverages additional computation and auxiliary cues to handle complex scenarios. Fast-slow systems enable flexible mode switching tailored to the varying demands of different scenarios. Chameleon (Zhang et al., 2025) utilizes a VLM with chain-of-thought (CoT) reasoning to conduct neuro-symbolic topology inference in the slow system. FAST (Xiao et al., 2025) dynamically switches between short and long reasoning paths based on task complexity. SlowFast-LLaVA (Xu et al., 2024) employs different sampling rates to capture motion cues and fuses slow and fast features to efficiently represent video information. Xu et al. (Xu et al., 2025) combine a “slow” LLM for command parsing with a “fast” RL agent for vehicle control. FASIONAD++ (Qian et al., 2025) incorporates a VLM as the slow system to provide feedback and evaluation for a fast end-to-end pipeline. However, in many of these approaches, the fast and slow systems are implemented as entirely separate frameworks, and switching occurs based on the scenario. In contrast, our method does not require additional models for the fast or slow pathways. Instead, fast and slow systems are trained under parallel supervision, enabling mutual reinforcement and improving overall robustness.
3 Method
3.1 Problem Formulation
@@
3.2 Overview
Fig. 2 illustrates the framework of FASTopoWM. The overall architecture can be briefly divided into encoder and decoder parts. Both the slow and fast systems are integrated into a unified network. The PV-to-BEV encoder (Li et al., 2022; He et al., 2016; Lin et al., 2017) extracts BEV features , where C, H, W represent the number of feature channels, height, and width, respectively. The initialized queries are derived from a learnable embedding space. is the number of queries. The relative ego poses between adjacent frames are encoded into an action latent . We employ memories for world models to store the historical queries. Conditioned on this action latent, the query world model and the BEV world model transform the historical queries and BEV features , respectively, to generate the stream queries and stream BEV features for the current frame. The stream BEV features are then fused with the BEV features extracted from the current frame. The first transformer layer takes the initialized queries and the BEV features from the current frame as input. The subsequent transformer layers share weights and receive parallel inputs from both the slow and fast systems. The slow system incorporates temporal information, whereas the fast system does not. During training, both systems are supervised jointly. At inference time, the prediction from the slow system is used as the final output when reliable pose information is available. Otherwise, when such information is missing or inaccurate, the prediction from the fast system is adopted. This design improves robustness while leveraging temporal cues to enhance performance when conditions permit.
3.3 Temporal Propagation via Latent World Models
Temporal propagation leverages detection results from previous frames as auxiliary information to assist prediction in the subsequent frame (Yuan et al., 2024). In contrast to existing latent world model methods (Li et al., 2024), our approach eliminates the need for future trajectory prediction. Instead, grounded in the principle that “the present represents a continuation of the past,” we utilize known relative poses to infer the current state based on past observations.
@@ 3.3 Temporal Propagation via Latent World Models
where query world model is composed of Transformer blocks, containing self-attention modules and feed-forward modules. Similarly, the BEV world model comprises Transformer blocks that incorporate temporal self-attention modules (Li et al., 2022) and feed-forward modules.
Then, the stream BEV feature are fused with extracted BEV feature using gated recurrent unit (Chung et al., 2014) to enrich temporal cues.
Future Latent Supervision. Unlike previous methods that rely on warping and often lose information at the BEV boundary (Wang et al., 2024b), our BEV world model imagines the evolution of the next-frame BEV representation based on historical BEV features and relative poses. Since dense BEV annotations are difficult to obtain, we adopt a self-supervised learning strategy based on BEV features from temporally adjacent frames. We use mean squared error (MSE) loss to align the stream BEV features with the extracted BEV features of the current frame:
For queries, we employ transformation loss (Yang et al., 2025b) to supervise consistency of coordinate, category, and semantic BEV mask:
where , , and are coordinates of lane segment, classes of centerline, boundary types and semantic BEV mask predicted from stream queries . , , and are GT annotations transformed from T-1 frame to T frame. For brevity, we omit the weights for each loss term. More details can be found in appendix. The overall future latent supervision can be expressed as:
@@
3.4 Unified Fast-Slow Decoder